Calculating the concentration of alcohol in terms of Molarity

The Concentration of alcohol (CH3CH2OH) in blood, called the "blood alcohol conctration" or BAC, is given in units of grams of alchol per100mL of blood. The legal definition of intoxication, in many states of the United States, is that the BAC is 0.08 or higher. What is the concentration of alcohol,in terms of molarity, in blood if the BAC is 0.08?
Answer #1
given maximum intoxication =(0.08 g of CH3CH2OH per 100 ml of blood)
molecular weight of CH3CH2OH=12+3+12+2+16+1=46 g

==> number of moles of CH3CH2OH in 0.08 g=0.08/46=0.00174 moles
molarity=(number of moles contained in 1 lit of solution)

==>molarity=(number of moles of solute)/(volume of solution in lit)
==>molarity=0.00174/0.1=0.0174 mol/lit

concentration in terms of molarity=0.0174 mol/lit
answered by: Miley
Answer #2

0.08 is in g/mL, so to get g you must multiply by the 100mL

  • (0.08g/mL)(100mL)=8g

Divide 8g by 46.07g/mol to get mol

  • (8g)/(46.07g/mol)=0.174mol

To get Molarity you must find mol/L, so now you must find L

  • 100mL/1000=0.1L

Now find mol/L

  • (0.174mol)/(0.1L)=1.74M

The answer is 1.74M